Grade 11Inorganic Chemistry

1.08 g of a metal oxide on being reduced with hydrogen gives 1 g of the metal.calculate the equivalent weight of the metal ?????????????????????

preet kaur
7 Years agoGrade 11

1 Answer

Mayank
ApprovedApproved Tutor Answer7 Years ago
here we use the concept that equivalents of metal oxide = equivalents of metal
eq.of metal oxide=1.08/(E+8)  where E=eq. weight of METAL   // note eq. weight of oxygen in metal oxide is 8
eq. of metal=1/E
now equating these we get 
1.08/(E+8)=1/E  which gives E=100 HENCE THE ANS.....